OpenStudy (anonymous):

Find equations of lines containg the three sides triangle abc A(00) B(12,6) C(18,0)

OpenStudy (anonymous):

in point slope form

OpenStudy (anonymous):

AB: y=(1/2)x AC: y=0 BC: y=-x+18

OpenStudy (anonymous):

thanks ok few more find midpoints of AB BC and AC graph and label the points x y and z respectively using black the graph is like this domain of x values -3 to 22 and range of y values -6 to 13

OpenStudy (anonymous):

sorry. i can't imagine.